Shift report 20030111 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 11 2003 - 08:07:27 EST

  • Next message: Apache: "Shift report 20030111 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Jan 11 08:07:27 2003, Shift: 00:00-08:00
    Supervisor: sjs, with: bn
    Log book#13 pages 48 - 50
    
    Shift Summary:
    ------------------------------------------------------------
    Single store from 1:30 until 7:30, when beam lost with quench condition.
    
    Runs 7309-7323.   Primarily worked on mrs trigger issue (not resolved) and h2 tdc spectra (also not resolved)
    
    ~50K Trig 6 (FS) (runs 7320 to 7323)
    
    FS set to 1/8 A field (8 deg.).  MRS at 3kG (90 deg.).
    
    We see a significant BG rate in the detectors and this seems to be driving the T1 and T2 anode currents quite high.  
    
    Problem:  No apparent H2 time spectra. The MRS trigger seems to fire primarily on background events. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Jan 11 2003 - 08:12:05 EST